find the sum of the first 19 terms of the arithmetic series in which a1=2 and d=4.
thanks i hope you can help!

Answer by venugopalramana(3286)

find the sum of the first 19 terms of the arithmetic series in which a1=2 and d=4.
thanks i hope you can help!
formula is
sn = (n/2){2a+(n-1)d}
n=19.............a=2........d=4.so
s19= (19/2){2*2+(19-1)4}=19*76/2=19*38=722